OpenStudy (anonymous):

if the radii of the circular ends of a conical bucket in the shape of frustum of a cone which is 45cm high are 28cm and 7cm.find the capacity of the bucket in litres

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Okay... the volume of a frustum, you know the formula? :)

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

What is the formula, then?

OpenStudy (anonymous):

v = pi/3 (r^2+R^2+rR)h

terenzreignz (terenzreignz):

Yep. So what's the volume, then? By the way, volume = capacity...

OpenStudy (anonymous):

= pi/3 (28^2+7^2+28*7)*45 = 15435 pi =15435(22/7) =48510cm^3

terenzreignz (terenzreignz):

Are you sure you want to use 22/7 as pi? Maybe you could use \[\large \pi = 3.141592654\]\[\huge \pi=3.1416\] Although if 22/7 is allowed, I won't contest it.

OpenStudy (anonymous):

because it is not given in question that u must use pi as 22/7 or 3.14

terenzreignz (terenzreignz):

Well, I don't want to assume. I'll just leave a factor of pi, and leave the deciding which value for pi you want to use, ok? :) So... \[\huge V=\frac{h\pi}{3}\left(R^2 + Rr + r^2\right)\]

OpenStudy (anonymous):

thank u so much @terenzreignz

terenzreignz (terenzreignz):

Not yet done! So your volume is \[\huge V=15435\pi\] But this is in cubic centimetres. You need the volume in litres...

OpenStudy (anonymous):

i have to divide it by 1000

terenzreignz (terenzreignz):

Okay, good :) It seems you already knew how to do it, why did you tag me? LOL Good job, by the way, but I'd suggest you use 3.1416 instead of 22/7 But anyway... nicely done

OpenStudy (anonymous):

thank u once again